OpenStudy (anonymous):

What value of r makes the equation true? r + 0.12 = 0.20 A. r = 0.32 B. r = 0.08 C. r = –0.08 D. r = –0.32

OpenStudy (anonymous):

Take a wild guess. XD What do you think it might be?

OpenStudy (anonymous):

Umm idk O_O thats why i asked XD

OpenStudy (mrnood):

don't make a wild guess. This is MATHS - it is a rational process work it out subtract 0.12 from each side of the equation

OpenStudy (anonymous):

@Cowboy17

OpenStudy (anonymous):

B @west_coast

OpenStudy (anonymous):

Thanks :)

OpenStudy (mrnood):

@west_coast you have no way of knowing if @paolaxx33 just guessed that. MANY answers given by people here are just plain WRONG

OpenStudy (anonymous):

Thanks i am thinking about it XD

OpenStudy (anonymous):

-0.08 i got that

OpenStudy (mrnood):

hmmm - well you need to try again r + 0.12 = 0.20 so subtract 0.12 from each side: r = 0.20- 0.12

OpenStudy (anonymous):

0.08?

OpenStudy (anonymous):

r + 0.12 = 0.20 -0.12 -0.12 r=0.08 To check you just add 0.08+0.12= 0.20 @west_coast

OpenStudy (mrnood):

that's better

OpenStudy (anonymous):

XD thanks

OpenStudy (anonymous):

so its 0.08

OpenStudy (anonymous):

yes @west_coast

OpenStudy (anonymous):

ok thanks